OpenStudy (anonymous):

Determine the vertex form of the equation of a parabola with focus (-2,-4) and directrix y=-6

OpenStudy (anonymous):

I've been working on this question for an hour and I can't figure it out..

OpenStudy (tkhunny):

Have you considered the definition? Parabola: Locus of Points Equidistant from a given point (focus) and a given line (directrix). What is the distance of any point from (-2,-4)? \(\sqrt{(x+2)^{2}+(y+4)^{2}}\) What is the distance of any point from y = -6 \(|y+6|\) Where are they the same? \(\sqrt{(x+2)^{2}+(y+4)^{2}} = |y+6|\) Everything is positive, so square both sides and use your best algebra.

OpenStudy (anonymous):

the vertex is half way between the focus and the directrix, can you find what this point would be?